High-power experimental engines are being developed by the Stevens Motor Company for use in its new sports coupe. The engineers have calculated the maximum horsepower for the engine to be 810HP. Nine engines are randomly selected for horsepower testing. The sample has an average maximum HP of 840 with a standard deviation of 50HP. Assume the population is normally distributed.

Step 1 of 2 :

Calculate a confidence interval for the average maximum HP for the experimental engine. Use a significance level of α=0.1. Round your answers to two decimal places.
Answer #1

df = n - 1 = 9 -1 = 8

From T table,

t critical value at 0.1 significance level with 8 df = 1.860

90% confidence interval for is

- t * S / sqrt(n) < < + t * S / sqrt(n)

840 - 1.86 * 50 / sqrt(9) < < 840 + 1.86 * 50 / sqrt(9)

809.00 < < 871.00

90% C I is ( 809.00 , 871.00 )